# Interaction as Explanation: A User Interaction-based Method for Explaining Image Classification Models

### v1

@@ 1 Introduction
Significant advancements have been achieved across multiple disciplines based on deep neural network models. Notably, models have evolved substantially in computer vision to address several tasks, such as image classification, object recognition, and semantic segmentation. From convolutional neural network (CNN) to Transformer-based models, the computer vision models have demonstrated high performance on various datasets and tasks [1, 8, 4]. The practical applications of these models are extensive and diverse, including autonomous driving, optical character recognition, medical diagnostic imaging, and commercial product recommendation.
@@
Despite the considerable advancements of deep neural networks, their deployment in the real world is fraught with challenges. Among these, the ’black-box’ problem is a particularly prominent challenge [22, 20]. This issue refers to the opacity and interpretability of deep neural network-based models, wherein the algorithms and operational principles lack transparency, making it difficult for users to comprehend the model’s decision-making process. The black-box problem can engender unpredictability in the outcomes generated by the models, thereby potentially affecting users’ trust [7]. Therefore, it is imperative to devise solutions addressing the black-box problem to apply models across diverse sectors successfully.
@@
Thus, prior research has proposed explainable AI (xAI) methods to mitigate the black-box problem within computer vision. The representative method is a feature-based explanation method, highlighting the input features that affect the models’ output and decision [3, 12, 10, 21, 17, 18, 13]. For instance, the Local Interpretable Model-Agnostic Explanations (LIME) technique identifies areas within input images that influence the outputs of image classification models [17]. Similarly, the gradient-weighted Class Activation Mapping (Grad-CAM) technique produces saliency maps to visualize important features based on gradients from the last layer of CNN models [18]. These techniques provide intuitive visual explanations for better model understanding. Other feature-based methods include the Testing with Concept Activation Vectors (TCAV) technique, partial dependence plots (PDP), and individual conditional expectation (ICE) methods. TCAV technique measures the impact of concepts from input images on the model’s output [10]. The partial dependence plots (PDP) and individual conditional expectation (ICE) methods illustrate how changes in inputs influence model outputs using a plot [3, 12, 9, 5]. These techniques are instrumental in highlighting the most significant features of the inputs. These xAI methods could be divided based on several criteria [12, 3, 11], such as the representation of explanations (e.g., heatmap, concepts, plots), model dependency (model-specific vs. model-agnostic), and the scope of explanation (global vs. local).
@@
The xAI methods play a crucial role by enabling users to comprehend how models process inputs [15, 3, 12, 16]. Furthermore, depending on the nature of the model and the specifics of the explanation needed, model designers could implement the appropriate xAI technique. However, previous investigations into xAI methods have primarily focused on explanations for AI experts [11]. While end-users also require explainability and interpretability on models, current xAI methods may not fully address their requirements [14]. The end-users want to participate in the design of xAI methods [11]; however, these methods are currently developed without incorporating end-user feedback.
@@
We propose an xAI method that enables users to understand models by interacting with image classification tasks. To facilitate the interaction-based xAI method, we have developed a web-based prototype. This prototype features a canvas that displays an image of a specific object, allowing users to paint over the image directly on the canvas. Initially, users examine the classification results of the image—including the class and confidence score— before it has been painted. Subsequently, they paint over areas that they believe do not influence the classification results. Then, the painted parts of the image are masked and removed from the image classification. After painting, they check how the classification results of the masked image differ from the initial classification results. By repeating this process, users can identify critical image features that influence the model’s input-to-output process and verify if their mental model is consistent with the model’s classification criteria. This approach can be universally applied to any classification model via our prototype.
@@ 2.1 Prototype for Interaction-based xAI Method
Our prototype was developed as a web-based system. We utilized the React library for frontend development and the FastAPI framework for backend development. The Pytorch was used for the deep learning framework [19]. Our prototype consists of a canvas with an image of an object, a section showing the classification results along with the confidence score, buttons for painting mode and erasing mode, and a button to check the classification results.
@@
The backend server runs the model’s inference on the image received from the frontend and responds to the request of the frontend. Thus, we first define an image classification model. The ResNet-50 model [6], pre-trained on the ImageNet dataset [2], is employed as the model. Then, the server preprocesses the image; it is resized and normalized according to the ImageNet dataset standards. After that, the model performs the inference on the image. Finally, the server applies the softmax function on the model’s output and sends the classification result –corresponding to the image class with the highest confidence– back to the frontend.
@@ 4 Conclusions
However, our work has several limitations. First, the method was applied solely to a basic image classification model. Given that computer vision models encompass a wide range of tasks, it is crucial to design interactions tailored to each specific task. Furthermore, our work focused on basic interactions, but exploring other interaction types, such as pinch, zoom, and tap, could be beneficial. Moreover, future considerations might include not only the creation of masked images but also modifications to the current image. Lastly, our research did not involve a user study; we only observed the changes in classification results from sample images. To ascertain whether there is an improvement in users’ understanding of computer vision models, conducting a user study is essential. Consequently, our future work aims to design a broader range of interactions, consider various computer vision tasks and models, and conduct a rigorous user study. Despite the limitations, our method is significant for proposing an xAI method that leverages end-users’ interactions.

### v2

@@ 1 Introduction
Significant advancements have been achieved across multiple disciplines based on deep neural network models. Notably, models have evolved substantially in computer vision to address several tasks, such as image classification, object recognition, and semantic segmentation. From convolutional neural network (CNN) to Transformer-based models, the computer vision models have demonstrated high performance on various datasets and tasks Arkin et al. (2021); Khan et al. (2022); Elngar et al. (2021). The practical applications of these models are extensive and diverse, including autonomous driving, optical character recognition, medical diagnostic imaging, and commercial product recommendation.
@@
Despite the considerable advancements of deep neural networks, their deployment in the real world is fraught with challenges. Among these, the ’black-box’ problem is a particularly prominent challenge Yun et al. (2023); Von Eschenbach (2021). This issue refers to the opacity and interpretability of deep neural network-based models, wherein the algorithms and operational principles lack transparency, making it difficult for users to comprehend the model’s decision-making process. The black-box problem can engender unpredictability in the outcomes generated by the models, thereby potentially affecting users’ trust Jiang et al. (2023). Therefore, it is imperative to devise solutions addressing the black-box problem to apply models across diverse sectors successfully.
@@
Thus, prior research has proposed explainable AI (xAI) methods to mitigate the black-box problem within computer vision. The representative method is a feature-based explanation method, highlighting the input features that affect the models’ output and decision Dwivedi et al. (2023); Liao and Varshney (2021); Kim et al. (2018); Yeh et al. (2018); Ribeiro et al. (2016); Selvaraju et al. (2017); Mothilal et al. (2020). For instance, the Local Interpretable Model-Agnostic Explanations (LIME) technique identifies areas within input images that influence the outputs of image classification models Ribeiro et al. (2016). Similarly, the gradient-weighted Class Activation Mapping (Grad-CAM) technique produces saliency maps to visualize important features based on gradients from the last layer of CNN models Selvaraju et al. (2017). These techniques provide intuitive visual explanations for better model understanding. Other feature-based methods include the Testing with Concept Activation Vectors (TCAV) technique, partial dependence plots (PDP), and individual conditional expectation (ICE) methods. TCAV technique measures the impact of concepts from input images on the model’s output Kim et al. (2018). The partial dependence plots (PDP) and individual conditional expectation (ICE) methods illustrate how changes in inputs influence model outputs using a plot Dwivedi et al. (2023); Liao and Varshney (2021); Khater et al. (2023); Goldstein et al. (2015). These techniques are instrumental in highlighting the most significant features of the inputs. These xAI methods could be divided based on several criteria Liao and Varshney (2021); Dwivedi et al. (2023); Kim et al. (2023), such as the representation of explanations (e.g., heatmap, concepts, plots), model dependency (model-specific vs. model-agnostic), and the scope of explanation (global vs. local).
@@
The xAI methods play a crucial role by enabling users to comprehend how models process inputs Patrício et al. (2023); Dwivedi et al. (2023); Liao and Varshney (2021); Recio-García et al. (2021). Furthermore, depending on the nature of the model and the specifics of the explanation needed, model designers could implement the appropriate xAI technique. However, previous investigations into xAI methods have primarily focused on explanations for AI experts Kim et al. (2023). While end-users also require explainability and interpretability on models, current xAI methods may not fully address their requirements Narwaria (2022). The end-users want to participate in the design of xAI methods Kim et al. (2023); however, these methods are currently developed without incorporating end-user feedback. Users frequently have xAI-related questions, such as ”Why did the system produce this result?”, ”Which features are significant?”, and ”Under what circumstances do the results change?” Liao et al. (2020). However, existing methods often struggle to provide intuitive answers to these questions. To address this challenge, previous research has proposed xAI interfaces that facilitate iterative user interaction Shneiderman (2020); Weld and Bansal (2019). These interfaces have been shown to enhance user satisfaction and comprehension of the system, while also facilitating active exploration Shneiderman (2020); Du et al. (2019).
We propose an interaction-based xAI method that enables users to understand models by observing changes in image classification results through painting-based interactions. This approach allows users to easily express their xAI questions by directly masking areas of interest. Furthermore, our approach enables users to immediately verify differences in classification results, thereby intuitively inferring answers to their questions. For instance, in the context of medical image-based diagnostic classification tasks, traditional methods typically represent the importance of individual image pixels through visualizations such as heat maps. However, these representations may not be readily comprehensible to patients or medical experts. In contrast, our interaction-based method allows experts or patients to mask suspected tumor regions and directly observe the resultant changes in diagnostic classification.
@@
To facilitate the interaction-based xAI method, we have developed a web-based prototype. This prototype features a canvas that displays an image of a specific object, allowing users to paint over the image directly on the canvas. Initially, users examine the classification results of the image—including the class and confidence score— before it has been painted. Subsequently, they paint over areas that they believe do not influence the classification results. Then, the painted parts of the image are masked and removed from the image classification. After painting, they check how the classification results of the masked image differ from the initial classification results. By repeating this process, users can identify critical image features that influence the model’s input-to-output process and verify if their mental model is consistent with the model’s classification criteria. This approach can be universally applied to any classification model via our prototype.
@@ 2.1 Prototype for Interaction-based xAI Method
Our prototype was developed as a web-based system. We utilized the React library for frontend development and the FastAPI framework for backend development. The Pytorch was used for the deep learning framework Steiner et al. (2019). Our prototype consists of a canvas with an image of an object, a section showing the classification results along with the confidence score, buttons for painting mode and erasing mode, and a button to check the classification results.
@@
The backend server runs the model’s inference on the image received from the frontend and responds to the request of the frontend. Thus, we first define an image classification model. The ResNet-50 model He et al. (2016), pre-trained on the ImageNet dataset Deng et al. (2009), is employed as the model. Then, the server preprocesses the image; it is resized and normalized according to the ImageNet dataset standards. After that, the model performs the inference on the image. Finally, the server applies the softmax function on the model’s output and sends the classification result –corresponding to the image class with the highest confidence– back to the frontend.
@@ 4 Conclusions
However, our work has several limitations. First, the method was applied solely to a basic image classification model. Given that computer vision models encompass a wide range of tasks, it is crucial to design interactions tailored to each specific task. Furthermore, our work focused on basic interactions, but exploring other interaction types, such as pinch, zoom, and tap, could be beneficial. Moreover, future considerations might include not only the creation of masked images but also modifications to the current image. Lastly, our research did not involve a user study; we only observed the changes in classification results from sample images. To ascertain whether there is an improvement in users’ understanding of computer vision models, conducting a user study is essential. Consequently, our future work aims to design a broader range of interactions, consider various computer vision tasks and models, and conduct a rigorous user study. Furthermore, we consider the approach that integrates our method with existing xAI techniques to provide a combination of automatic explanations and additional interactions, potentially reducing the challenges associated with iteration in the interaction-based method. Despite the limitations, our method is significant for proposing an xAI method that leverages end-users’ interactions.